Please help

1. Directions for reconstitution of mezlocillin of IV injection: "For IV use reconstitute with 30mL sterile water for injection. Each 30 mL of reconstituted solution will contain 3 grams of mezlocillin." The client is to receive 1.5 grams of mezlocillin IV every 6 hours. How many milliliters would the nurse administer to the client?

2. Directions for reconstitution of ampicillin sodium for injection: "For IM use add 3.5 mL sterile water for injection. Resulting solution contains 250 mg ampicillin sodium per mL. The vial contains ampicillin sodium equivalent to 1 gram. Usual adult dosage of ampicillin is 250 mg–500 mg IM every 6 hours." How much diluent is added to the vial to prepare themedication for IM use?

3. A child weighs 18 kg and has an order for Zithromax 10 mg/kg po stat. The recommended dosage range on day one is 10 mg/kg. The oral suspension is dispensed as 200 mg in 5 mL. How many milliliters will you administer in one dose?

4. The client is to receive Xanax (alprazolam) 0.5 mg po @hs. What time would be appropriate to administer this medication?

5. A child weighs 40 kg and the recommended dosage of Cleocin (clindamycin palmitate hydrochloride) oral suspension is 8 to 25 mg/kg/day in four divided dosages. What is the maximum dosage for this child in 24 hours?

6. Directions for reconstitution of oxacillin sodium for injection: "For IM use add 5.7 mL sterile water for injection. Resulting solution contains 250 mg oxacillin sodium per 1.5 mL. The vial contains oxacillin sodium equivalent to 1 gram. Usual adult dosage of oxacillin sodium is 250 mg–500 mg IM every 4 to 6 hours." The client is to receive 400 mg IM every 6 hours. How many milliliters would the nurse administer to the client?
Answer #1

1.Given

Route:IV

Frequency:every six hour

Available 30ml=3g (Mezlocillin)

Order:1.5gm

When 3gm=30ml

Then 1.5g=??ml

Cross multiply and divide

=30×1.5÷3

=15

Ans:15ml

2.Given

Diluent:Sterile water

Amount:3.5ml is diluted to get 250mg of ampicillin sodium per ml

So here the quantity of diluent used to get the medication is 3.5ml

Ans:3.5ml diluent is added to the vial to prepare the medication

3.Given

Medication:Zithromax

Route:oral

Frequency:at once

Weight:18kg

Order:10mg/kg

So for 18 kg multiply with 10

=180kg(18×10)

The disease is 180mg

Available:200mg=5ml

So 180mg=??ml

Cross multiply and divide

=180×5÷200

=900÷200

=4.5

Ans:4.5ml has to be administered in one dose

4.Given

Order:Xanax

Dose:0.5mg

Route:oral

Frequency: HS

The HS stands for Hora somni ,in simple the patient supposed to take this medicine at bed time

Ans: It is best to administer the medication at bed time

5.Given

Order:Cleocin

Route:Oral

Frequency:four times

Dose:8_25mg/kg/day

Weight:40kg

So for 8 kg cross multiply and divide (40×8÷1)=320mg

For 25mg/kg cross multiply and divide (25×40÷1)=1000mg

The recommended dose is 320mg to 1000mg per day

Ans:The maximum dosage for the child in 24 hours should be 1000mg